OpenStudy (6l4d3):

Use the Newton-Raphson iteration procedure to find the root of 𝑥^3−6𝑥^2+9𝑥+1 = 0

ganeshie8 (ganeshie8):

Hi

OpenStudy (mww):

You will need at least one approximation to the root(s) to get you started. Differentiating we get 3x^2 - 12x + 9 = 3(x^2 - 4x + 3) = 3(x-3)(x-1) = 0 so x = 3 and x = 1 are our relative maxima and minima. P(3) = 1 and P(1) = 5 So we anticipate it will look something like this:  So expect the root to occur for x < 0. But sub in points and see what x gets the cubic closer to 0. If your x becomes more negative but the cubic becomes more negative, we need to head in the opposite direction. Once you find a suitable 'alpha', then apply Newton's law to it.